# Envelopes and evolutes

### v1

@@ 4. Evolutes of curves
Let $f:X\to\mathbb{P}(V)$ be a morphism from a nonsingular curve $X$. Assume $f(X)$ is in general position with respect to $H_{\infty}$ and $Q_{\infty}$. Let $d=f^{*}c_{1}(\mathcal{O}_{\mathbb{P}(V)}(1))\cap[X]$ denote the degree of $X$, $g$ its genus, and $k_{0}$ the degree of the branch locus of $f$, i.e., $k_{0}$ is the (weighted) number of cusps of $f(X)$. Let $\mathcal{P}^{1}$ denote the image of the first jet map $V_{X}\to\mathcal{P}^{1}_{X}(1)$, and set $\mathcal{K}^{1}:=\operatorname{Ker}V_{X}\to\mathcal{P}^{1}$ (cf. Remark 3.1).
@@
When $n=2$, so that $X$ is a plane curve, the envelope is the evolute of the curve. See [ev] for a thorough treatment of this case, both in the real and complex situation.
@@
Consider now the case $n=3$. Then the cuspidal edge of the envelope is the evolute of $X$. We note that the formulas for the degrees of the envelope and of the evolute are in accordance with Salmon’s formulas [Salmon]*Footnote, p. 341. He states that the envelope $E_{\mathcal{E}}$ has degree $3m+n$, where $m$ is the degree of the curve $X$ and $n$ is the degree of its strict dual $X^{*}$. In the notation of [numchar], we have $m=d$ and $n=r_{2}=3(d+2g-2)-2k_{0}-k_{1}$, hence $3m+n=3d+3(d+2g-2)-2k_{0}-k_{1}=6(d+g-1)-2k_{0}-k_{1}$. Salmon assumes that $X$ has no inflection points, i.e., that $k_{1}=0$, so in this case his formula agrees with the one of Proposition 4.1. In a footnote he says that if $X$ has $\theta$ inflection points, then $n$ should be replaced by $n+\theta$, so his formula would read $3m+n+\theta=3d+3(d+2g-2)-2k_{0}-k_{1}+k_{1}=6(d+g-1)-2k_{0}$, in agreement with ours.
@@
Salmon’s formula for the degree of the evolute $C_{\mathcal{E}}$ is $5m+\alpha$, where $\alpha$ denotes the number $k_{2}=4(d+3(g-1))-3k_{0}-2k_{1}$ of hyperosculating planes to $X$, i.e., $5m+\alpha=5d+4(d+3(g-1))-3k_{0}-2k_{1}=3(3d+4g-4-k_{0})-2k_{1}$. This agrees with our formula in Proposition 4.3 when $k_{1}=0$. But the number of hyperosculating points should be weighted, so that $\alpha=2k_{1}+k_{2}$. This gives $5m+\alpha=5d+4(d+3(g-1))-3k_{0}-2k_{1}+2k_{1}=3(3d+4g-4-k_{0})$, again in agreement with the formula in Proposition 4.3.
To find the degree of the evolute $C_{\mathcal{E}}$, Salmon [Salmon]*Footnote, p. 341 asserts that the degree of the curve of osculating planes to the evolute is $m+r$, where $m$ is the degree of $X$ and $r$ is the rank of $X$, i.e., the degree of its tangent developable. The reason for this is that the osculating planes to the evolute are the normal planes to $X$.
@@ 5. Evolutes of surfaces
Assume $\mathbb{P}(V)$ has a Euclidean structure as described in the beginning of Section 4. Let $X\subset\mathbb{P}(V)$ be a nonsingular surface which is in general position with respect to $H_{\infty}$ and $Q_{\infty}$ [reci]*Prop. 4.1, p. 245. Then there is a natural surjective map $V_{X}\to\mathcal{E}$, where $\mathcal{E}=(\mathcal{K}^{1})^{\vee}\oplus\mathcal{O}_{X}(1)$ is the Euclidean normal bundle of $X$. Note that $\mathcal{K}^{1}=\mathcal{N}(1)$ is the twisted conormal bundle of $X$ in $\mathbb{P}(V)$. Then $\psi:\mathbb{P}(\mathcal{E})\to\mathbb{P}(V)$ is the family of normal (n2)(n-2)-spaces to $X$.
@@
We apply Corollaries 2.5 and 2.9 with $\mathcal{F}=\mathcal{E}$. We have
$c_{1}(\mathcal{E})=-c_{1}(\mathcal{K}^{1})+c_{1}(\mathcal{O}_{X}(1))=c_{1}(\mathcal{P}^{1}_{X}(1))+c_{1}(\mathcal{O}_{X}(1))=c_{1}(\Omega_{X}^{1})+4c_{1}(\mathcal{O}_{X}(1))$
and
@@
For $d=2$, our formula gives $\deg E_{\mathcal{E}}=12$, in accordance with Salmon [Salmon]*Art. 206, p. 179, who actually computes its equation. Moreover, in the fifth edition of Salmon’s book [Salmon2]*Art. 511, p. 151, he gives the formula for $\deg E_{\mathcal{E}}$ above and also the formula $2d(d^{2}-d-1)$ for the degree of the dual surface – the class of $E_{\mathcal{E}}$ [Salmon2]*Art. 509, p. 150. For $d=2$, this gives 4, as is also stated in [Salmon]*Art. 199, p. 171. In the general case, to find the class of the evolute, Salmon starts by determining the number of the bitangents of the evolute, which are the same as the normals to the surface $X$, that pass through a given point [Salmon2]*Art. 507, p. 149. This number is the socalled Euclidean distance degree of $X$ [Draisma]*, and is equal to $d+d(d-1)+d(d-1)^{2}=d(d^{2}-d+1)$. He shows this by taking the point to be a point in the plane at infiinity. From this, and a consideration of the points on a surface, the normals at which meet a given line, he deduces the formula for the class of $E_{\mathcal{E}}$. For the degree of $E_{\mathcal{E}}$, he gives two alternative approaches. One, attributed to Darboux, uses the theory of congruences of lines, the lines being the bitangents of the evolute, which are also the normals of the surface $X$. The other consists in studying the intersection of $E_{\mathcal{E}}$ with the plane at infinity.
@@
Porteous observed that the points on the evolute above the umbilics (where the two focal points coincide) are points of (Lagrangian) singularity type $D_{4}$ for the map $\psi:\mathbb{P}(\mathcal{E})\to\mathbb{P}(V)$ [P1]*p. 160. Hence their number should be computable as the value of the corresponding Thom polynomial $\widetilde{Q}_{21}=\tilde{c}_{1}\tilde{c}_{2}-2\tilde{c}_{3}$, where the $\tilde{c}_{i}$ are the Chern classes of the virtual bundle $\Omega_{\mathbb{P}(\mathcal{E})}^{1}-\psi^{*}\Omega_{\mathbb{P}(V)}^{1}$ [MPW]*Rmk. 13, p. 76. Unfortunately the computation does not give the desired formula.
Another approach, as suggested in [P2]*p. 124, for an affine surface $X_{0}\subset\mathbb{A}^{3}$, would be to consider the set of coinciding focal points as the $\Sigma^{2,2}$ locus of the map
$X_{0}\times\mathbb{A}^{3}\to\mathbb{A}^{1}\times\mathbb{A}^{3},$
@@ 6. Osculating developables of curves
Assume $X$ is a curve. Denote by $\mathcal{P}^{m}_{X}(1)$ the sheaf of principal parts of order $m$ of $\mathcal{O}_{X}(1):=f^{*}\mathcal{O}_{\mathbb{P}(V)}(1)$, and let $a^{m}\colon V_{X}\to\mathcal{P}^{m}_{X}(1)$ denote the $m$th jet map. Let $\mathcal{P}^{m}:=\operatorname{Im}a^{m}$ denote the $m$th osculating bundle [numchar]*Prop. (2.1), p. 478, and set $\mathcal{K}^{m}:=\operatorname{Ker}a^{m}$, so that we have exact sequences
@@
Note that
@@
This gives [numchar]*Thm. (3.2), p. 481
$\deg D^{m}_{X}=\binom{m+1}{2}(2g-2)+(m+1)d-\sum_{i=0}^{m-1}(m-i)k_{i},$
where $g$ is the genus of $C$ and $d$ its degree, and $k_{i}$ its $i$th stationary index (see [numchar]).
It is well know that in the case $n=2$, so that $X$ is a plane curve, the envelope of the tangents to $X$ is equal to $X$ union its inflectional tangents. When $n=3$, the envelope of the osculating planes to $X$ is equal to the tangent developable of $X$ union its hyperosculating planes. More generally, we have the following result.
@@
to $X$ is equal to its (n2)(n-2)th osculating developable $D^{n-2}_{X}$ union its hyperosculating hyperplanes. We have
$\deg E_{\mathcal{P}^{n-1}}=2\bigl{(}nd+(n^{2}-n+1)(g-1)-\sum_{i=0}^{n-2}(n-1-i)k_{i}\bigr{)}.$
@@
Let $r(t)$ be a local parameterization of $f\colon X\to\mathbb{P}(V)$. The map $\psi$ is given locally by
@@
Its differential is given by
$d\psi=\left(\begin{array}[]{c}r^{\prime}(t)+u_{1}r^{\prime\prime}(t)+\cdots+u_{n-1}r^{(n)}(t)\\ r^{\prime}(t)\\ \vdots\\ r^{(n-1)}(t)\end{array}\right)\cong\left(\begin{array}[]{c}u_{n-1}r^{(n)}(t)\\ r^{\prime}(t)\\ \vdots\\ r^{(n-1)}(t)\end{array}\right).$
Hence the rank of $d\psi$ is <n<n when $u_{n-1}=0$, i.e., when the point lies in the (n2)(n-2)th osculating space to $X$ at $r(t)$, or when $r^{(n)}(t)=0$, i.e., when the point $r(t)$ is a point of hyperosculation. ∎
It follows from Proposition 2.4 that
$\deg E_{\mathcal{P}^{n-1}}$ $=$ $2g-2+2c_{1}(\mathcal{P}^{n-1})\cap[X]$
$=$ $2g-2+2n\bigl{(}d+(n-1)(g-1)-\sum_{i=0}^{n-2}(n-1-i)k_{i}\bigr{)},$
and we note that the numerical formulas
$\deg D_{X}^{n-2}$ $=$ $(n-1)(d+(n-2)(g-1))-\sum_{i=0}^{n-3}(n-2-i)k_{i}$
$k_{n-1}$ $=$ $(n+1)(d+n(g-1))-\sum_{i=0}^{n-2}(n-i)k_{i}$
verify $\deg E_{\mathcal{P}^{n-1}}=\deg D_{X}^{n-2}+k_{n-1}$.
Example 6.2.
Let $f\colon\mathbb{P}^{1}\to\mathbb{P}^{n}$ be a rational normal curve. Then $f^{*}\mathcal{O}_{\mathbb{P}^{n}}(1)=\mathcal{O}_{\mathbb{P}^{1}}(n)$ and the family of osculating hyperplanes to $f(\mathbb{P}^{1})$ is given by
$\mathbb{P}(\mathcal{P}^{n-1}_{\mathbb{P}^{1}}(n))=\mathbb{P}(\mathcal{O}_{\mathbb{P}^{1}}(1)^{n})\subset\mathbb{P}^{1}\times\mathbb{P}^{n}.$
Since a rational normal curve has no points of hyperosculation, the envelope of its family of osculating hyperplanes is equal to its (n2)(n-2)th osculating developable $D_{f(\mathbb{P}^{1})}^{n-2}$, which has degree 2(n1)2(n-1).
Recall that the dual variety $X^{\vee}\subset\mathbb{P}(V^{\vee})$ of $X$ is defined as the image of the composed map
$\mathbb{P}((\mathcal{K}^{1})^{\vee})\subset X\times\mathbb{P}(V^{\vee})\to\mathbb{P}(V^{\vee}).$
The degree of $X^{\vee}$ is
$\deg X^{\vee}=c_{1}(\mathcal{O}_{\mathbb{P}((\mathcal{K}^{1})^{\vee}})(1))^{2}\cap[\mathbb{P}((\mathcal{K}^{1})^{\vee})]=c_{1}((\mathcal{K}^{1})^{\vee})\cap[X].$
Since $c_{1}((\mathcal{K}^{1})^{\vee})=c_{1}(\mathcal{P}^{1})$, we see that $\deg X^{\vee}=\deg D^{1}_{X}$.
The strict dual curve $X^{*}$ of $X$ is the image of the composed map
$\mathbb{P}((\mathcal{K}^{n-1})^{\vee})\subset X\times\mathbb{P}(V^{\vee})\to\mathbb{P}(V^{\vee}).$
Recall that $X^{**}=X$ holds, that the dual variety $X^{\vee}$ is equal to the (n1)(n-1)th osculating developable $D^{n-1}_{X^{*}}$ of $X^{*}$, and, vice versa, $X^{*\vee}=D^{n-1}_{X}$, see [numchar]. For a treatment of the case of real curves, see [Ishikawa].

### v2

@@ 4. Evolutes of curves
Let $f:X\to\mathbb{P}(V)$ be a morphism from a nonsingular curve $X$, and assume $f$ is birational onto its image. Assume $f(X)$ is in general position with respect to $H_{\infty}$ and $Q_{\infty}$ (see Remark 3.1). Let $d=f^{*}c_{1}(\mathcal{O}_{\mathbb{P}(V)}(1))\cap[X]$ denote the degree of $X$, and $g$ its genus.
Let $k_{0}$ denote the degree of the ramification locus of $f$, i.e., $k_{0}$ is the “weighted number of cusps” of $f(X)$. Since $X$ is a nonsingular curve, the image $\mathcal{P}^{1}$ of the first jet map $V_{X}\to\mathcal{P}^{1}_{X}(1)$, is a vector bundle, of rank 2. The cokernel of the jet map is equal to $\Omega^{1}_{X/\mathbb{P}(V)}(1)$. Hence we have an exact sequence
$0\to\mathcal{P}^{1}\to\mathcal{P}^{1}_{X}(1)\to\Omega^{1}_{X/\mathbb{P}(V)}(1)\to 0.$
The ramification locus is defined by the 0th Fitting ideal of $\Omega^{1}_{X/\mathbb{P}(V)}$, equivalently of $\Omega^{1}_{X/\mathbb{P}(V)}(1)$ [numchar]*p. 481. Hence we get
$c_{1}(\mathcal{P}^{1})\cap[X]=c_{1}(\mathcal{P}^{1}_{X}(1))\cap[X]-k_{0}=2d+2g-2-k_{0},$
where the last equality follows from the fact that $c_{1}(\mathcal{P}^{1}_{X}(1))=c_{1}(\Omega^{1}_{X})+2c_{1}(\mathcal{O}_{X}(1))$. Set $\mathcal{K}^{1}:=\operatorname{Ker}(V_{X}\to\mathcal{P}^{1})$. Then $\mathcal{E}=\mathcal{(}\mathcal{K}^{1})^{\vee}\oplus\mathcal{O}_{X}(1)$ is the Euclidean normal bundle (see Remark 3.1).
@@
When $n=2$, so that $f(X)$ is a plane curve, the envelope is the evolute of the curve. See [ev] for a thorough treatment of this case, both in the real and complex situation.
@@
Consider now the case $n=3$. Then the cuspidal edge $C_{\mathcal{E}}$ of the envelope $E_{\mathcal{E}}$ is the evolute of $X$. We shall see that the formulas for the degrees of the envelope and the evolute are in accordance with Salmon’s formulas [Salmon]*Footnote, p. 341. He states that the envelope $E_{\mathcal{E}}$ has degree $3m+n$, where $m$ is the degree of the curve $X$ and $n$ is the number of osculating planes to the curve that pass through a given point. Let $k_{1}$ denote the (weighted) number of inflection points of the curve. We have $m=d$ and $n=3(d+2g-2)-2k_{0}-k_{1}$ [numchar]*Thm. 3.2, p. 481. Hence $3m+n=3d+3(d+2g-2)-2k_{0}-k_{1}=6(d+g-1)-2k_{0}-k_{1}$. Salmon assumes that the curve has no inflection points, i.e., that $k_{1}=0$, so in this case his formula agrees with the one of Proposition 4.1. In a footnote he says that if the curve has $\theta$ inflection points, then $n$ should be replaced by $n+\theta$, so his formula would read $3m+n+\theta=3d+3(d+2g-2)-2k_{0}-k_{1}+k_{1}=6(d+g-1)-2k_{0}$, in agreement with Proposition 4.1
@@
Salmon’s formula for the degree of the evolute $C_{\mathcal{E}}$ is $5m+\alpha$, where $\alpha$ denotes the number $k_{2}=4(d+3(g-1))-3k_{0}-2k_{1}$ of hyperosculating planes to the curve [numchar]*Thm. 3.2, p. 481, i.e., $5m+\alpha=5d+4(d+3(g-1))-3k_{0}-2k_{1}=3(3d+4g-4-k_{0})-2k_{1}$. This agrees with our formula in Proposition 4.3 when $k_{1}=0$. But the number of hyperosculating points should be weighted, so that $\alpha=2k_{1}+k_{2}$. This gives $5m+\alpha=5d+4(d+3(g-1))-3k_{0}-2k_{1}+2k_{1}=3(3d+4g-4-k_{0})$, again in agreement with the formula in Proposition 4.3.
To find the degree of the evolute $C_{\mathcal{E}}$, Salmon [Salmon]*Footnote, p. 341 asserts that the degree of the curve in the dual projective space $\mathbb{P}(V^{\vee})$ consisting of the osculating planes to the evolute is $m+r$, where $m$ is the degree of $X$ and $r$ is the rank of $X$, i.e., the degree of its tangent developable. The reason for this is that the osculating planes to the evolute are the normal planes to $X$ [BL].
@@ 5. Evolutes of surfaces
Assume $\mathbb{P}(V)$ has a Euclidean structure as described in the beginning of Section 3, and that $X\subset\mathbb{P}(V)$ is a nonsingular surface which is in general position with respect to $H_{\infty}$ and $Q_{\infty}$. Then the Euclidean normal bundle is $\mathcal{E}=(\mathcal{K}^{1})^{\vee}\oplus\mathcal{O}_{X}(1)$, where $\mathcal{K}^{1}=\operatorname{Ker}(V_{X}\to\mathcal{P}^{1}_{X}(1))$ is equal to the conormal bundle of $X$ in $\mathbb{P}(V)$ twisted by $\mathcal{O}_{X}(1)$: The composed map $\psi:\mathbb{P}(\mathcal{E})\subset X\times\mathbb{P}(V)\to\mathbb{P}(V)$ is the family of normal (n2)(n-2)-spaces to $X$.
@@
We apply Corollaries 2.6 and 2.10 with $\mathcal{F}=\mathcal{E}=(\mathcal{K}^{1})^{\vee}\oplus\mathcal{O}_{X}(1)$. We have
$c_{1}(\mathcal{E})=-c_{1}(\mathcal{K}^{1})+c_{1}(\mathcal{O}_{X}(1))=c_{1}(\mathcal{P}^{1}_{X}(1))+c_{1}(\mathcal{O}_{X}(1))=c_{1}(\Omega_{X}^{1})+4c_{1}(\mathcal{O}_{X}(1)).$
Now $c(\mathcal{K}^{1})=c(\mathcal{P}^{1}_{X}(1))^{-1}$, which gives
$c_{2}(\mathcal{K}^{1})=c_{1}(\Omega^{1}_{X})^{2}-c_{2}(\Omega_{X}^{1})+4c_{1}(\Omega_{X}^{1})c_{1}(\mathcal{O}_{X}(1))+6c_{1}(\Omega_{X}^{1})^{2}.$
Hence we get
@@
For $d=2$, our formula gives $\deg E_{\mathcal{E}}=12$, in accordance with Salmon [Salmon]*Art. 206, p. 179, who actually computes its equation. Moreover, in the fifth edition of Salmon’s book [Salmon2]*Art. 511, p. 151, he gives the formula for $\deg E_{\mathcal{E}}$ above and also the formula $2d(d^{2}-d-1)$ for the degree of the dual surface – the class of $E_{\mathcal{E}}$ [Salmon2]*Art. 509, p. 150. For $d=2$, this gives 4, as is also stated in [Salmon]*Art. 199, p. 171. In the general case, to find the class of the evolute, Salmon starts by determining the number of the bitangents of the evolute, which are the same as the normals to the surface $X$, that pass through a given point [Salmon2]*Art. 507, p. 149. This number is the socalled Euclidean distance degree of $X$ [Draisma]*, and is equal to $d+d(d-1)+d(d-1)^{2}=d(d^{2}-d+1)$. He shows this by taking the point to be a point in the plane at infinity. From this, and a consideration of the points on a surface, the normals at which meet a given line, he deduces the formula for the class of $E_{\mathcal{E}}$. For the degree of $E_{\mathcal{E}}$, he gives two alternative approaches. One, attributed to Darboux, uses the theory of congruences of lines, the lines being the bitangents of the evolute, which are also the normals of the surface $X$. The other consists in studying the intersection of $E_{\mathcal{E}}$ with the plane at infinity.
@@
Porteous considered affine surfaces in $\mathbb{R}^{3}$ and observed that the points on the evolute above the umbilics (where the two focal points coincide) are points of singularity type $D_{4}$ for the map from the total space of the normals to the surface to $\mathbb{R}^{3}$. One could hope that this could carry over to the present situation of a complex projective surface, and that the map $\psi\colon\mathbb{P}(\mathcal{E})\to\mathbb{P}(V)$ would be Lagrangian. If so, the number of umbilics could be computed by evaluating the the corresponding Thom polynomial $\widetilde{Q}_{21}=\tilde{c}_{1}\tilde{c}_{2}-2\tilde{c}_{3}$, where the $\tilde{c}_{i}$ are the Chern classes of the virtual bundle $\Omega_{\mathbb{P}(\mathcal{E})}^{1}-\psi^{*}\Omega_{\mathbb{P}(V)}^{1}$ [MPW]* p. 76. However, this computation does not give Salmon’s formula, and the discrepancy is rather large.
Another approach, as suggested in [P2]*p. 124, again for an affine surface $X_{0}\subset\mathbb{R}^{3}$, would be to consider the set of coinciding focal points as the $\Sigma^{2,2}$ locus of the map
$X_{0}\times\mathbb{R}^{3}\to\mathbb{R}\times\mathbb{R}^{3},$
@@ 6. Osculating developables of curves
Assume $X$ is a nonsingular curve and $f\colon X\to\mathbb{P}(V)$ a map which is birational onto its image. Let $d=f^{*}c_{1}(\mathcal{O}_{\mathbb{P}(V)}(1))\cap[X]$ denote the degree of $X$, and $g$ its genus. Denote by $\mathcal{P}^{m}_{X}(1)$ the sheaf of principal parts of order $m$ of $\mathcal{O}_{X}(1):=f^{*}\mathcal{O}_{\mathbb{P}(V)}(1)$, and let $a^{m}\colon V_{X}\to\mathcal{P}^{m}_{X}(1)$ denote the $m$th jet map. Let $\mathcal{P}^{m}:=\operatorname{Im}a^{m}$ denote the $m$th osculating bundle [numchar]*Prop. (2.1), p. 478, and set $\mathcal{K}^{m}:=\operatorname{Ker}a^{m}$, so that we have exact sequences
@@
From the exact sequences
$0\to S^{i}\Omega^{1}_{X}\otimes\mathcal{O}_{X}(1)=(\Omega^{1}_{X})^{\otimes i}\otimes\mathcal{O}_{X}(1)\to\mathcal{P}^{i}_{X}(1)\to\mathcal{P}^{i-1}_{X}(1)\to 0,$
it follows that
@@
The stationary indices $k_{i}$ of the curve are defined pointwise via local parameterizations of the curve [numchar]*p. 481. Intuitively, $k_{0}$ is the (weighted) number of cusps, $k_{1}$ the (weighted) number of inflectional tangents, …, and $k_{n-1}$ the (weighted) number of hyperosculating hyperplanes. These numbers satisfy the following global formulas [numchar]*Thm. (3.2), p. 481:
$(c_{1}(\mathcal{P}^{m}_{X}(1))-c_{1}(\mathcal{P}^{m}))\cap[X]=\sum_{i=0}^{m-1}(m-i)k_{i}.$
Hence
$\deg D^{m}_{X}=c_{1}(\mathcal{P}^{m})\cap[X]=\binom{m+1}{2}(2g-2)+(m+1)d-\sum_{i=0}^{m-1}(m-i)k_{i}.$
It is well know that in the case $n=2$, when $X$ is a plane curve, the envelope of the tangents to $X$ is equal to $X$ union its inflectional tangents. When $n=3$, the envelope of the osculating planes to $X$ is equal to the tangent developable of $X$ union its hyperosculating planes. More generally, we have the following result, which generalizes and extends to the complex case a similar result of Ishikawa for real curves [Ishikawa]*Thm. 1, p. 604.
@@
to $X$ is equal to its (n2)(n-2)th osculating developable $D^{n-2}_{X}$ union its hyperosculating hyperplanes.
@@
Let $t$ be a local parameter on $X$ and $t,u_{1},\dots,u_{n-1}$ local parameters on $\mathbb{P}(\mathcal{P}^{n-1})$. Let $r(t)$ be a local parameterization of $f\colon X\to\mathbb{P}(V)$. The (n1)(n-1)th jet map $a^{n-1}\colon V_{X}\to\mathcal{P}^{1}_{X}(1)$ is given locally by the matrix
$A^{n-1}(t):=\left(\begin{array}[]{c}r(t)\\ r^{\prime}(t)\\ \vdots\\ r^{(n-1)}(t)\end{array}\right).$
Above points on $X$ where the map $a^{n-1}$ is surjective, the map $\psi\colon\mathbb{P}(\mathcal{P}^{n-1})\to\mathbb{P}(V)$ is given locally by
@@
Its differential $d\psi$ is given by
$dA^{n-1}(t,u):=\left(\begin{array}[]{c}r^{\prime}(t)+u_{1}r^{\prime\prime}(t)+\cdots+u_{n-1}r^{(n)}(t)\\ r^{\prime}(t)\\ \vdots\\ r^{(n-1)}(t)\end{array}\right)\cong\left(\begin{array}[]{c}u_{n-1}r^{(n)}(t)\\ r^{\prime}(t)\\ \vdots\\ r^{(n-1)}(t)\end{array}\right).$
By assumption, the rank of $A^{n-1}(t)$ is $n$, hence the rank of $d\psi$ is <n<n when $u_{n-1}r^{(n)}(t)=0$. If $u_{n-1}=0$, the point lies in the (n2)(n-2)th osculating space to $X$ at $r(t)$, hence in $D^{n-2}_{X}$. If $r^{(n)}(t)=0$, the point $r(t)$ is a point of hyperosculation, and $d\psi$ has rank <n<n at all points above it in $\mathbb{P}(\mathcal{P}^{n-1})$, i.e., at all points in the (n1)(n-1)th osculating space at the point $r(t)$.
In the case that we are considering a point $r(t)$ where $a^{n-1}$ is not surjective, say for $t=0$, then the inclusion $\mathcal{P}^{n-1}\subseteq\mathcal{P}^{n-1}_{X}(1)$ is given by a matrix which has rank <n<n for $t=0$. The map $V_{X}\to\mathcal{P}^{n-1}$ is given by a modification of the matrix $A^{n-1}(t)$. A local study as in [numchar]*pp.478–479 shows that we get the same conclusion as in the previous case. ∎
Let us check that the numerical formulas agree. It follows from Proposition 2.5 that
$\deg E_{\mathcal{P}^{n-1}}=2g-2+2c_{1}(\mathcal{P}^{n-1})\cap[X].$
hence
$\deg E_{\mathcal{P}^{n-1}}=2g-2+2n\bigl(d+(n-1)(g-1)-\sum_{i=0}^{n-2}(n-1-i)k_{i}\bigr)\\ \textstyle=2\bigl(nd+(n^{2}-n+1)(g-1)-\sum_{i=0}^{n-2}(n-1-i)k_{i}\bigr).$
From what we have seen,
$\deg D_{X}^{n-2}=(n-1)(d+(n-2)(g-1))-\sum_{i=0}^{n-3}(n-2-i)k_{i}.$
We also know that [numchar]*Thm. (3.2), p. 481
$k_{n-1}=(n+1)(d+n(g-1))-\sum_{i=0}^{n-2}(n-i)k_{i}.$
It follows that $\deg E_{\mathcal{P}^{n-1}}=\deg D_{X}^{n-2}+k_{n-1}$.